OpenStudy (octoberroo):

PLZ HELP FOR MEDALS Solve x2 - 3x = -8

pooja195 (pooja195):

Are we supposed to use the quadratic formula?

pooja195 (pooja195):

Or factoring...? What do the directions say?

OpenStudy (anonymous):

what are your options

pooja195 (pooja195):

@EllenJaz17 its better to guide the user rather than do the work for them

OpenStudy (octoberroo):

3+-i sqrt29 x= --------- 2 3+-i sqrt23 x= --------- 2 -3+-i sqrt29 x= --------- 2 -3+-i sqrt23 x= --------- 2

OpenStudy (octoberroo):

those are the choices can someone please help

pooja195 (pooja195):

Are you fimilar with this equation? \[\huge~\rm~ax^2+bx+c=0\]

OpenStudy (anonymous):

start with adding 3 to both sides

OpenStudy (octoberroo):

i just want to know the answer honestly

OpenStudy (octoberroo):

thank you<3

pooja195 (pooja195):

Congrats for not learning anything.

OpenStudy (solomonzelman):

when the equation is \(\LARGE\color{black}{ \color{blue}{a} x^2+ \color{magenta}{b}x+ \color{red}{c}=0 }\). then, x solutions go by the following formula: \(\bbox[8pt, lightyellow ,border:8px solid black]{\huge \huge{x=~} \huge{ \frac{-\color{magenta}{b} \pm\sqrt{ \color{magenta}{b} ^2-4 \color{blue}{a} \color{red}{c}}}{2 \color{blue}{a}} } ~ }\)

OpenStudy (solomonzelman):

you just need to add 8 to both sides, and to identify your abc.